# Semantic Information Management in Low-Temperature Plasma Science and Technology with VIVO

### v1

@@ Abstract
Digital research data management is increasingly integrated across universities and research institutions, addressing the handling of research data throughout its lifecycle according to the FAIR data principles (Findable, Accessible, Interoperable, Reusable). Recent emphasis on the semantic and interlinking aspects of research data, e.g., by using ontologies and knowledge graphs further enhances findability and reusability. This work presents a framework for creating and maintaining a knowledge graph specifically for low-temperature plasma (LTP) science and technology. The framework leverages a domain-specific ontology called Plasma-O, along with the VIVO software as a platform for semantic information management in LTP research. While some research fields are already prepared to use ontologies and knowledge graphs for information management, their application in LTP research is nascent. This work aims to bridge this gap by providing a framework that not only improves research data management but also fosters community participation in building the domain-specific ontology and knowledge graph based on the published materials. The results may also support other research fields in the practical use of knowledge graphs for semantic information management.
@@ 1 Introduction
The assimilation of digital research data management (RDM) is seeing a growing trend across universities and research institutes. This is visible from the increasing adoption rate of electronic laboratory notebooks (ELNs) and other RDM software tools [Petersen et al. (2020), Fink et al. (2023), Schröder et al. (2022), Chaerony Siffa et al. (2022)]. This trend focuses not only on the technical aspects of data handling throughout the research lifecycle but also emphasizes the importance of semantic information management in enhancing the findability, understandability, and reusability of research data. RDM practices primarily focus on the technical aspects of data handling, ensuring the availability and accessibility of research data through collection, storage, archiving, and publication on institutional or domain-specific repositories [Conrad et al. (2024), Lin et al. (2024), Rahimzadeh et al. (2023)]. However, semantic information management goes beyond organizing and storing data by adding layers of meaning and context through the use ontologies, knowledge graphs, and other semantic technologies [Schneider and Šimkus (2020), Schröder et al. (2022)]. This semantic enrichment of research data unlocks its full potential for knowledge discovery and reuse, enabling researchers to identify relationships between datasets, understand the context in which data was collected, and assess its quality and relevance for specific research questions. By making research data machine-readable and understandable, semantic information management enhances the interoperability and reusability of data, aligning with the FAIR (Findable, Accessible, Interoperable, Reusable) data principles [Wilkinson et al. (2016)].
@@
In the field of low-temperature plasma (LTP) science and technology, often characterized by multi-disciplinary research topics and how experiments are conducted (small table-top experiments as opposed to large-scale experiments as common, e.g. in high-energy physics) [Anirudh et al. (2023)], the interlinking and explicit relationships between relevant concepts can be practical. For example, a researcher may be interested in finding datasets that are obtained from a certain diagnostic method applied to study a certain plasma source. This can be modeled by assigning a relationship between the concepts of plasma source and dataset, i.e. it can be described in an ontology, which is a formal description of concepts or knowledge and their relationships within a domain. In this example, the use of an ontology further improves the findability and comprehensibility of complex data. The work on ontologies applied in the LTP field is still in its infancy. To the best of our knowledge, there has only been a couple of academic articles on ontology development in the field of computational plasma physics [Snytnikov et al. (2020), Sapetina et al. (2020)], which is adjacent to the LTP field. Notably, the first steps in that direction have been taken within the framework of developing a metadata schema and the foundation of a general knowledge graph for applied plasma physics and plasma medicine [Sack and Hoppe (2023), Becker et al. (2023), Becker et al. (2020)].
@@ 2 Methods
This section details the development and integration of a domain-specific ontology tailored for the field of low-temperature plasma (LTP) science and technology. The overarching objective is to establish a robust semantic framework that facilitates knowledge creation, discovery, and management within the LTP domain. This is achieved through two primary approaches: the construction of a specialized ontology, Plasma-O, and its subsequent integration into the VIVO software to create and maintain the knowledge graph.
@@ 2.1 Ontology for Low-Temperature Plasma Science and Technology
The domain-specific ontology presented in this work is built upon the earlier draft [Sack and Hoppe (2023), Becker et al. (2023), Becker et al. (2020)] and the concepts introduced in the existing plasma metadata schema, named Plasma-MDS, introduced by Franke et al. [Franke et al. (2020)]. There, the introduced metadata schema is intended to support the adoption of the FAIR data principles and RDM in the LTP field, specifically for the publication of datasets generated from LTP experiments. Plasma-MDS is used as the metadata standard for LTP-specific data repositories such as INPTDAT [^1] https://www.inptdat.de/ [Becker et al. (2019)] and RDPCIDAT [^2] https://rdpcidat.rub.de/. The developed ontology, however, is designed to have a broader application range, e.g. as a basis of metadata schemas for various RDM workflows and semantic knowledge creation and discovery. In this work, we focus on the part of the domain-specific ontology for the latter case and consider the first case and the complete ontology for future work. From here on, we shall refer to this domain-specific ontology as the Plasma Ontology or Plasma-O.
In contrast with Plasma-MDS that introduces Dataset as a central class or concept, Plasma-O is built around the central class PlasmaStudy to which other classes are connected. It is formally defined (not a full definition) as
$PlasmaStudy$ $\sqsubseteq\geq 1hasResearchTopic.ResearchTopic$
$\sqsubseteq\geq 1isConductedBy.\text{foaf:}Agent$
$\sqsubseteq\exists hasOutput.\left(\text{bibo:}Document\sqcup\text{vivo:}Dataset\right)$
$\sqsubseteq\exists usesDevice.DiagnosticDevice$
$\sqsubseteq\exists usesMethod.DiagnosticMethod$
@@
This design choice is to ensure the applicability of the ontology outside the dataset publication use case, and establishes Dataset as an output of PlasmaStudy as per the designated relationship.
Another core class in Plasma-O is Device, a generic class encompassing both PlasmaSource and DiagnosticDevice. The key distinction lies in the specific relationships PlasmaSource has with other classes connected to PlasmaStudy. For instance, DiagnosticDevice lacks direct relationships with Target and Medium, which are exclusive to PlasmaSource. It is important to note that Plasma-O, by design, includes inverse properties for most object properties. This means that if PlasmaStudy connects to PlasmaSource through the hasSource property, the inverse relationship can be expressed using the isSourceOf property.
@@
Figure 1: Selected Plasma-O classes and their relationships with each other and classes from external ontologies.
@@
Plasma-O can be attached to other ontologies by means of reusing existing classes from them. Here, we attach Plasma-O to the VIVO ontology [Corson-Rikert et al. (2012)], which forms the backbone of the VIVO software. The VIVO ontology itself reuses classes and properties from other ontologies such as FOAF [Brickley and Miller (2014)], BIBO [D’Arcus and Giasson (2009)], and OBO [Smith et al. (2007)], to name a few. Together, these ontologies are used to describe various different concepts that can be grouped into people, activities, courses, events, organizations, equipment, research, and locations. This encompasses the general concepts in most scientific disciplines and highlights the interoperability of the different concepts from these ontologies. Figure 1 illustrates an example of selected classes from Plasma-O attached to the VIVO ontology (which includes the classes from OBO, FOAF, and BIBO ontologies) and their basic relations. Note that classes and properties that are not preceded by a namespace belong to Plasma-O.
@@
Plasma-O, despite being able to describe general relevant concepts in LTP, is still being developed further. The presented work strives to provide public access to the ontology and related knowledge graph going—with respect to accessibility for humans and machines—far beyond storing the source files in a public repository. The suggested approach of using VIVO as a semantic knowledge management system for domain-specific data and metadata (beyond its original application scenario) is seen as transferable to other domains. With this, it potentially fosters the adoption of the FAIR data principles in different interdisciplinary research communities. The version of Plasma-O (version 0.5.0) used in the present work is available at the Plasma-MDS GitHub page [^3] https://github.com/plasma-mds/plasma-ontology.
@@ 2.2 LTP Knowledge Graph Development with VIVO
The core technology of VIVO is the ontologies. VIVO provides a user-friendly user interface (UI) that enables end users (e.g. data curators) to create and manage semantic information in the form of a knowledge graph based on the available ontologies. As mentioned in Section 2.1, VIVO comes with the VIVO ontology and various other ontologies readily facilitating the creation of a knowledge graph representing researchers, their interests, projects, and outputs. Additionally, incorporating new ontologies and/or knowledge graphs into VIVO is simplified through its add/remove RDF data feature. Accordingly, we introduce Plasma-O into VIVO and set the namespace (or prefix) to “plasma-o” for quick identification of the ownership of the classes. Note that, we use VIVO version 1.13 for this work. For a more detailed documentation to install and set up VIVO version 1.13, we refer the readers to the relevant VIVO documentation [^4] https://wiki.lyrasis.org/display/VIVODOC113x.
@@
There are several ways to populate the ontologies with instances or data on VIVO to create a knowledge graph ab intio. The most straightforward way is to use the data input feature. As a first step, we create a class group, called “plasma-o”, and pick relevant Plasma-O classes (and also relevant classes from other related ontologies) to be included in the plasma-o group. This is done to ease the data ingest process on VIVO. We can then choose the starting class we want to populate with instances. Figure 2 shows an example workflow for creating an instance of PlasmaSource. It starts from selecting the PlasmaSource class (available under the plasma-o group). Then we will be prompted to provide a name for this instance, in this case, we choose the technical name for this plasma source, which is “kINPen-sci”. Upon creating the instance, we will be directed to the newly created kINPen-sci instance page as shown in step 3 in Figure 2. Since PlasmaSource is connected to different other classes in the ontology through the corresponding properties, these connections will be automatically displayed on the kINPen-sci instance page. Here, we can proceed to link the kINPen-sci instance with either existing instances or new instances. In the example, we link the kINPen-sci instance with an already existing PlasmaStudy instance, through the isSourceOf property (shown by its label “is plasma source of” in the figure). A similar process can be done for other classes and properties. In addition to creating the instance one by one, which can be time consuming, VIVO also allows for creation of instances and linking in bulk using CSV files. This particular approach streamlines the process by allowing users to provide instance details like id, name, uri, and other relevant properties in a structured format.
@@ 3 Results
The primary and most general use case of VIVO is to create profiles for researchers and their research outputs, typically in the form of published academic articles and other contributions. This directly relates to the expertise these researchers possess. For instance, a researcher might be linked to several publications and listed as experienced in specific plasma sources, devices, or diagnostic methods. Beyond this, we demonstrate other potential use cases of VIVO within the LTP field, which can be adapted by other domains as well. Some of these potential use cases are demonstrated in this section.

### v3

@@ Abstract
Digital research data management is increasingly integrated across universities and research institutions, addressing the handling of research data throughout its lifecycle according to the FAIR data principles (Findable, Accessible, Interoperable, Reusable). Recent emphasis on the semantic and interlinking aspects of research data, e.g. by using ontologies and knowledge graphs further enhances findability and reusability. This work presents a framework for creating and maintaining a knowledge graph specifically for low-temperature plasma (LTP) science and technology. The framework leverages a domain-specific ontology called Plasma-O, along with the VIVO software as a platform for semantic information management in LTP research. While some research fields are already prepared to use ontologies and knowledge graphs for information management, their application in LTP research is nascent. This work aims to bridge this gap by providing a framework that not only improves research data management but also fosters community participation in building the domain-specific ontology and knowledge graph based on the published materials. The results may also support other research fields in the practical use of knowledge graphs for semantic information management.
@@ 1 Introduction
The assimilation of digital research data management (RDM) is seeing a growing trend across universities and research institutes. This is visible from the increasing adoption rate of electronic laboratory notebooks (ELNs) and other RDM software tools [Petersen et al. (2020), Fink et al. (2023), Schröder et al. (2022), Chaerony Siffa et al. (2022), Jackson et al. (2024)]. This trend focuses not only on the technical aspects of data handling throughout the research lifecycle but also emphasizes the importance of semantic information management in enhancing the findability, understandability, and reusability of research data. RDM practices primarily focus on the technical aspects of data handling, ensuring the availability and accessibility of research data through collection, storage, archiving, and publication on institutional or domain-specific repositories [Conrad et al. (2024), Lin et al. (2024), Rahimzadeh et al. (2023), Strand et al. (2022)]. However, semantic information management goes beyond organizing and storing data by adding layers of meaning and context through the use ontologies, knowledge graphs, and other semantic technologies [Schneider and Šimkus (2020), Schröder et al. (2022)]. This semantic enrichment of research data unlocks its full potential for knowledge discovery and reuse, enabling researchers to identify relationships between datasets, understand the context in which data was collected, and assess its quality and relevance for specific research questions. By making research data machine-readable and understandable, semantic information management enhances the interoperability and reusability of data, aligning with the FAIR (Findable, Accessible, Interoperable, Reusable) data principles [Wilkinson et al. (2016)].
@@
In the field of low-temperature plasma (LTP) science and technology, often characterized by multi-disciplinary research topics and how experiments are conducted (small table-top experiments as opposed to large-scale experiments as common, e.g. in high-energy physics) [Anirudh et al. (2023)], the interlinking and explicit relationships between relevant concepts can be practical. For example, a researcher may be interested in finding datasets that are obtained from a certain diagnostic device used to diagnose a certain plasma source. This can be modeled by assigning a relationship between the concepts of plasma source and dataset, i.e. it can be described in an ontology, which is a formal description of concepts or knowledge and their relationships within a domain. In this example, the use of an ontology further improves the findability and comprehensibility of complex data. To illustrate further, Figure 1a portrays a setup of a table-top experiment typically used in LTP research, which comprises several interconnected instances (components). Correspondingly, Figure 1b provides a prototypical schematic representation of the relationships between these components, highlighting the interplay between the plasma source, the treated sample, and the diagnostic devices participating in the experiment, which ultimately results in the generation of research data.
The work on ontologies applied in the LTP field is still in its infancy. To the best of our knowledge, there has only been a couple of academic articles on ontology development in the field of computational plasma physics [Snytnikov et al. (2020), Sapetina et al. (2020)], which is adjacent to the LTP field. Notably, the first steps in that direction have been taken within the framework of developing a metadata schema and the foundation of a general knowledge graph for applied plasma physics and plasma medicine [Sack and Hoppe (2023), Becker et al. (2023), Becker et al. (2020)].
Figure 1: (a) Typical experiment setup found in low-temperature plasma research. (b) Schematic representation of the relationships between the components within the experiment.
@@ 2 Methods
This section details the development and integration of a domain-specific ontology tailored for the field of low-temperature plasma (LTP) science and technology. The overarching objective is to establish a robust semantic framework that facilitates knowledge creation, discovery, and management within the LTP domain. This is achieved through two primary approaches: the construction of a specialized ontology, Plasma-O, and its subsequent integration into the VIVO platform to create and maintain the knowledge graph.
@@ 2.1 Ontology for Low-Temperature Plasma Science and Technology
The domain-specific ontology presented in this work is built upon the earlier draft [Sack and Hoppe (2023), Becker et al. (2023), Becker et al. (2020)] and the concepts introduced in the existing plasma metadata schema, named Plasma-MDS, introduced by Franke et al. [Franke et al. (2020)]. There, the introduced metadata schema is intended to support the adoption of the FAIR data principles and RDM in the LTP field, specifically for the publication of datasets generated from LTP experiments. Plasma-MDS is used as the metadata standard for LTP-specific data repositories such as INPTDAT [^1] https://www.inptdat.de/ [Becker et al. (2019)] and RDPCIDAT [^2] https://rdpcidat.rub.de/. The developed ontology in this work, however, is designed to have a broader application range, e.g. as a basis of metadata schemas for various RDM workflows and semantic knowledge creation and discovery. In this work, we focus on the part of the domain-specific ontology for the latter case and consider the first case and the complete ontology for future work. From here on, we shall refer to this domain-specific ontology as the Plasma Ontology or Plasma-O.
While Plasma-MDS serves effectively as a metadata schema for describing and publishing LTP research datasets, Plasma-O, as a formal ontology encompasses a broader scope. Many of the terminologies and concepts present in Plasma-MDS are also defined within Plasma-O, but crucially, Plasma-O extends beyond data description. It establishes formal relationships between these concepts, enabling automated reasoning, inference of new knowledge, and integration of data from various sources. This richer semantic representation allows Plasma-O to support a wider range of RDM workflows, including data analysis, modeling, and complex semantic queries. This helps in facilitating advanced knowledge discovery and turning the FAIR data principles into practice. In essence, Plasma-O is complementary to Plasma-MDS, adding the possibility of semantic linking of data and metadata.
In contrast to Plasma-MDS that introduces Dataset as a central class or concept, Plasma-O is built around the central classes Plasma Study and Plasma Experiment, which are special cases and subclasses of Scientific Study and Scientific Experiment, respectively. Both may share a similar set of relationships with other classes that makes it natural to group them within the same superclass, i.e. Research Activity. Formal expressions can be introduced to ensure the specificity of Plasma Study and Plasma Experiment, which set them apart from the generic Scientific Study and Scientific Experiment. For example, the formal expression for Plasma Experiment may be written as
$Plasma\,Experiment$ $\sqsubseteq\geq 1involves.Plasma$
$\sqsubseteq\exists conductedBy.Agent$
$\sqsubseteq\exists hasTopic.Research\,Topic$
@@
$\sqsubseteq\exists usesMethod.Method$
$\sqsubseteq\exists hasOutput.Document$
$\sqsubseteq\exists usesDevice.Device$
The first property (or relation) of the expression above denotes that every Plasma Experiment must involve at least one kind of Plasma. The remaining properties relate Plasma Experiment to other classes, such as Plasma Source, Medium, Target, Device, etc. This is consistent with the concepts introduced in Plasma-MDS [Franke et al. (2020)]. Referring to Figure 1b, these classes or concepts indeed generalize over various components in the experiment, e.g. the gas $\mathrm{N_{2}}$ and $\mathrm{O_{2}}$ can be classified as Medium, water as Target, gliding arc plasma source as Plasma Source (which is modeled as a subclass of Device), etc. From the given expression, one can infer the experiment illustrated in Figure 1 is of type Plasma Experiment considering it involves an arc discharge, which is an instance of Plasma. Furthermore, the output of an experiment is Document, which may be in form of Dataset and/or Article, amongst others. Subsequently, one can write an expression of Plasma Study as concisely as follows
$Plasma\,Study$ $\sqsubseteq\left[\left(\geq 1consistsOf.Plasma\,Experiment\right)\sqcup\left(\geq 1involves.Plasma\right)\right]$
which means that every Plasma Study must have at least one Plasma Experiment or involve at least a kind of Plasma. Although in practice, this class also shares a similar set of properties like Plasma Experiment, which are inherited from the superclass Research Activity.
[…]
It is worth noting that Plasma-O, by design, includes inverses for most object properties. This means that if Research Activity is connected to Medium through the hasMedium property, the inverse relationship can be expressed using the mediumOf property. These inverse properties come in handy when exploring the classes and their instances in Plasma-O and later in the constructed knowledge graph. Figure 2 illustrates a simplified representation of Plasma-O consisting of selected classes and their relationships relevant for the manuscript. Furthermore, Table 1 provides descriptions of the classes.
@@
Figure 2: Simplified excerpt of Plasma-O illustrating selected classes and their relationships with each other and classes from external ontologies. Inverse properties are not visualized for simplicity.
Table 1: Descriptions of selected classes from Plasma-O and reused classes from other ontologies.
Class Ontology Description
Research Activity Plasma-O Any systematic investigation or inquiry aimed at increasing knowledge or understanding.
Plasma Experiment Plasma-O A scientific experiment that involves plasma as the object of investigation and/or its use in various kinds of applications, e.g. plasma medicine, decontamination, material processing, etc.
Plasma Study Plasma-O A collection of experiments with the focus on investigating plasma behaviors/properties, effects of a plasma on a certain target, its applications, etc.
Device Plasma-O A device is an object that is invented or constructed to serve a specific function. Also equivalent to vivo:Instrument.
Plasma Source Plasma-O A device that generates a plasma. It is a subclass of Device
Medium Plasma-O A medium such as gas, mixture of gases, or other substance, which is either (partially) ionized or otherwise activated by a plasma source.
Target Plasma-O An entity, which a plasma acts on.
Method Plasma-O Equivalent to obo:Technique: a technique is a planned process used to accomplish a specific activity or task.
Configuration Plasma-O A set of settings and parameters of a device or instrument.
Agent FOAF Agents are things that do stuff. Subclasses of this class include Person, Institute, Group, etc.
Document BIBO A document (noun) is a bounded physical representation of body of information designed with the capacity (and usually intent) to communicate. A document may manifest symbolic, diagrammatic or sensory-representational information.
Dataset VIVO Dataset is a document consisting of a collection of digital resources obtained from scientific experiments. It is a subclass of Dataset.
@@
Plasma-O can be attached to other ontologies by means of reusing existing classes from them. Here, we attach Plasma-O to the VIVO ontology [Corson-Rikert et al. (2012)], which forms the backbone of the VIVO platform. The VIVO ontology itself reuses classes and properties from other ontologies such as FOAF [Brickley and Miller (2014)], BIBO [D’Arcus and Giasson (2009)], and OBO [Smith et al. (2007)], to name a few. Together, these ontologies are used to describe various different concepts that can be grouped into people, activities, courses, events, organizations, equipment, research, and locations. This encompasses the general concepts in most scientific disciplines and highlights the interoperability of the different concepts from these ontologies. Some concrete examples include the reuse of Agent from FOAF connected to Plasma Experiment via the conductedBy property, and Document from BIBO via hasOutput. Apart from reuse, Plasma-O also considers class equivalences, e.g. Method is equivalent to Technique from OBO, and Device is equivalent to Instrument from VIVO.
@@
Despite being able to describe general relevant concepts in LTP research, Plasma-O is still being developed further. The presented work strives to provide public access to the ontology and related knowledge graph going—with respect to accessibility for humans and machines—far beyond storing the source files in a public repository. The suggested approach of using VIVO as a semantic knowledge management system for domain-specific data and metadata (beyond its original application scenario) is seen as transferable to other domains. With this, it potentially fosters the adoption of the FAIR data principles in different interdisciplinary research communities. The version of Plasma-O introduced in this work (version 0.7.0) [^3] Release version 0.7.0 is available at https://github.com/plasma-mds/plasma-ontology/releases/tag/v0.7.0 consists of 61 classes, 105 object properties, and 12 individuals whose numbers are expected to grow in the next iteration of the ontology.
@@ 2.2 LTP Knowledge Graph Development with VIVO
The core technology of VIVO is the ontologies. VIVO provides a user-friendly user interface (UI) that enables end users (e.g. data curators) to create and manage semantic information in the form of a knowledge graph based on the available ontologies. As mentioned in Section 2.1, VIVO comes with the VIVO ontology and various other ontologies readily facilitating the creation of a knowledge graph representing researchers, their interests, projects, and outputs. Additionally, incorporating new ontologies and/or knowledge graphs into VIVO is simplified through its add/remove RDF data feature. Given the aforementioned features and existing resources in VIVO, we see the opportunity to extend VIVO’s original use case that is current research information system with a more detailed semantic knowledge management system for research objects in LTP research. Compared to other systems such as Semantic MediaWiki [^4] https://www.semantic-mediawiki.org and Wikibase [^5] https://wikiba.se/, VIVO offers out-of-the-box flexibility and simplicity for content integration and reuse for both humans and machines, which we see as a big advantage.
Accordingly, we introduce Plasma-O into VIVO and set the namespace (or prefix) to “plsmo” for quick identification of the ownership of the classes. Note that, we use VIVO version 1.15 for this work. For a more detailed documentation to install and set up VIVO version 1.15, we refer the readers to the relevant VIVO documentation [^6] https://wiki.lyrasis.org/display/VIVODOC115x.
@@
There are several ways to populate the ontologies with instances or data on VIVO to create a knowledge graph ab intio. The most straightforward way is to use the data input feature. As a first step, we create a class group, called “plsmo”, and pick relevant Plasma-O classes (and also relevant classes from other related ontologies) to be included in the plsmo group. This is done to ease the data ingest process on VIVO. We can then choose the starting class we want to populate with instances. Figure 3 shows an example workflow for creating an instance of Plasma Source. It starts from selecting the Plasma Source class (available under the plasma-o group). Then we will be prompted to provide a name for this instance, in this case, we choose the technical name for this plasma source, which is “kINPen-sci”. Upon creating the instance, we will be directed to the newly created kINPen-sci instance page as shown in step 3 in Figure 3. Since Plasma Source is connected to different other classes in the ontology through the corresponding properties, these connections will be automatically displayed on the kINPen-sci instance page. Here, we can proceed to link the kINPen-sci instance with either existing instances or new instances. In the example, we link the kINPen-sci instance with an already existing Plasma Study instance, through the plasmaSourceOf property. A similar process can be done for other classes and properties. In addition to creating the instance one by one, which can be time consuming, VIVO also allows for creation of instances and linking in bulk using CSV files. This particular approach streamlines the process by allowing users to provide instance details like id, name, uri, and other relevant properties in a structured format.
@@ 3 Results
The primary and most general use case of VIVO is to create profiles for researchers and their research outputs, typically in the form of published academic articles and other contributions. This directly relates to the expertise these researchers possess. For instance, a researcher might be linked to several publications and listed as experienced in specific plasma sources, devices, or diagnostic methods. Beyond this, we demonstrate other potential use cases of VIVO within the LTP field, which can be adapted by other domains as well.